OpenStudy (anonymous):

Help! Please! I will give medal! Find the distance between the two points using the distance formula: Point C: 3,2 Point D: 7,-2

OpenStudy (anonymous):

Do you know about the distance formula?

OpenStudy (anonymous):

yeah but every time I get 0

OpenStudy (anonymous):

\[d = \sqrt{( x _{2}- x _{1})^{2} + (y _{2}-y _{1})^{2}}\] Set your problem up. \[d = \sqrt{(7 - 3)^{2}+(-2 + 2)^{2}}\] Simplify. \[d=\sqrt{(4)^{2} + (0)^{2}}\] That 0 gets erased and you're left with... \[\sqrt{16}\]

OpenStudy (anonymous):

And so your distance will be 4.

OpenStudy (anonymous):

oh my gosh thank you so much!

OpenStudy (anonymous):

You're welcome :)